OpenStudy (anonymous):

For which equation would x = 4 be a solution? 4x + 7 = 23 4x ÷ 8 = 8 8 - 5x = 1 7 + 3x = 18

OpenStudy (anonymous):

4x+7=23 4(4)+7=23 16+7=23 23=23